Homework #8 Due Monday, November 18 1. Decide whether a precipitate will form when the following solutions are mixed (use Tab part B please
Answer #1

Solubility equillibrium of CaF2 is

CaF2(s) <-------> Ca2+(aq) + 2F-(aq)

Ksp = [ Ca2+] [ F-]2= 3.9 × 10-11

Qsp = [Ca2+]initial × [F-]2initial

Total volume = 1.0ml + 10.0ml + 100ml = 111ml

[Ca2+]initial = 0.0030M/(111.0ml / 10ml)) = 0.0002703M

[F-]initial = 1.0M/ ( 111.0ml/ 1.0ml) = 0.009009M

Qsp = 0.0002703M × ( 0.009009M)2 = 2.19 × 10-8

Qsp > Ksp

Therefore

Precipitate will form